OpenStudy (anonymous):

What are the first four terms of the sequence defined by t1 = 15 and tn = tn-1 + 3.

OpenStudy (anonymous):

so T1 = 15. so T2 = T1 + 3 = 15+3 = 18 T3 = 21, Hence T4 = 24.